1,2,2,4,4,4,8,8,8,8,.........
find the 1000th term.

Grade:10

1 Answers

Yash Sinha
12 years ago

20= taken 1 times

21= taken 2 times

.

.

.

2n= taken n+1 times

hence total terms = 1+2+3+...+n+1=(n+1)(n+2)/2= 1000

                                                 or,n2+3n+2=2000

                                                 or,n2+3n-1998=0

                                                     n=(-3+√9+4*1998)/2= 43.22~43

Hence 1000th term is 243

 

 

Please  acknowledge the correct result